The age of a father is 5 more than twice of his son's age. The age of the mother is the sum of her son's age and one third of the father's age. The sum of all their ages is 100. Find their ages.

Answer by TimothyLamb(4379)

f = 2s + 5
m = s + f/3
m + f + s = 100
---
m + f + s = 100
s + f/3 + 2s + 5 + s = 100
s + (2s + 5)/3 + 2s + 5 + s = 100
s + (2/3)s + 5/3 + 2s + 5 + s = 100
4s + (2/3)s = 100 - 5/3 - 5
(12/3)s + (2/3)s = 300/3 - 5/3 - 15/3
(14/3)s = 280/3
14s = 280
---
s = 20
f = 45
m = 35
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php